john's father is now twice as old as he is. 8 years ago, he was thrice as old. How old is each now?
Answer by Maths68(1474)

Let
Present age of father = f
Present age of John = j
john's father is now twice as old as he is
f=2j ..........(1)
8 years ago means we have to subtract 8 from their present ages
f-8=j-8
then, he was thrice as old.
f-8=3(j-8)
f-8=3j-24
f=3j-24+8
f=3j-16...........(2)
Put the value of f from (1) to (2)
f=3j-16...........(2)
2j=3j-16
2j-3j=-16
-j=-16
Multiply by -1 both sides of the equation
-1*-j=-1*-16
j=16
Put the value of j in (1)
f=2j ..........(1)
f=2(16)
f=32
Present age of father = f = 32 years old
Present age of John = j = 16 years old